AZ Lucifer run - Night #7

Observers: Dave, Marie, Fuyan

TO: Geno

Scripts observed saved on lucifer in: /home/observer/AZ/May2011/

Data prefix:20110506_

15:00 --Running darks from last night

#1 -- 50 normal mode darks

#51 -- 80 integrated mode darks -- Telescope started slewing during frame 50. We will redo these integrated darks tomorrow.

19:45 Significant cloud cover. Running Fuyan's bootes imaging script which does not require photometric conditions.

Seeing ~1.4

#81 -- 95

20:40 Seeing ~1" going to acquire A1703 ( mos_a1703_acq.txt)

#96 -- 100 acquisition frames

Seeing ~0.6

20:55 Starting h-band spectroscopy of a1703 (mos_a1703_sci_h.txt)

#101 -- 124

Seeing ~0.5

23:10 Going to telluric acquisition for a1703 (mos_tell_a1703_acq.txt)

#126 --129 acquisition frames

23:20 Doing H-band telluric measurements (mos_tell_a1703_sci_h.txt)

#130 -- 133 H-band spectra

23:30 Running h-band flats for a1703 (mos_a1703_flt_h.txt)

#134 -- 139

23:35 Seeing 0.5 going to telluric for gn26 (gn26_1.0arcsec/gn26_calstar_acq.txt )

Problem with pointing going to do pointing check

23:50 Thin cloud cover moving in (< .5 mag extinction). Redoing telluric acquisition for gn26 (gn26_1.0arcsec/gn26_calstar_acq.txt )

#141 -- 143 acquisition frames

23:55 Doing telluric spectra (gn26_1.0arcsec/gn26_calstar_exp.txt)

#144 -- 145 spectra

24:00 Going to gn26 science acquisition (gn26_1.0arcsec/gn26_main_acc.txt)

#146 -- 148 acquisition frames

Seeing 0.8", thin cloud cover

00:10 Starting science script for gn26 (gn26_1.0arcsec/gn26_main_exp.txt)

#149 --154 stable seeing < 1", thin cloud cover increasing with time during script

00:45 Running flats for gn26 (gn26_1.0arcsec/gn26_ff.txt)

#156 --160

Seeing 0.7"

00:50 Going to acquire bootes telluric (bootes_tel.acq.txt)

#161 -image of telluric star

Went to put mask in FPU and had a MOS error. "Error setting the MOS position" --looks like mask did not reach limit switch.

Dave is debugging with Walter.

1:30 This is a problem that the MPE folks need to solve. Trying to get in touch with them.

2:00 Peter is working on the problem.

2:15 Peter has fixed the MOS unit. Correcting for pointing.

2:25 Seeing ~0.8". We will only have 1.5 hours of science left tonight (accounting for acquisitions and tellurics) so we will not do the bootes field tonight.

Starting acquisition script for a1703 (mos_a1703_acq.txt)

#163 -- 167 acquisition frames

2:40 Starting K-band spectral observations of a1703 (mos_a1703_sci_k.txt)

#168 -- 185 Seeing steadily worse, ~1.7--2 at the end.

4:20 Going to telluric for a1703 (mos_a1703_acq.txt)

#186 -- 190 acquisition frames

4:30 K spectra of telluric (mos_tell_a1703_sci_k.txt)

#191 -- 194

4:40 K-band flats for a1703 ( mos_a1703_flt_k.txt)

#195 -- 200

4:45 Running end of night script

5:00 Running darks

-- XiaohuiFan - 06 May 2011
Topic revision: r2 - 06 May 2011, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback